5. Distribute an amount of Rs. 200 between Raheem and Usman
such that Raheem gets Rs.50 more than twice as much as Usman
gets?

8. Distribute an amount of Rs.500 among 2 brothers and 1 sister
such that,
a. sister gets twice as much as brothers do.
b. each brother gets twice as much as the sister does.

Answer by macston(5194)

5. Distribute an amount of Rs. 200 between Raheem and Usman
such that Raheem gets Rs.50 more than twice as much as Usman
gets?
.
R=amount to Raheem; U=amount to Usman
R=2U+50 . Use this to substitute for R below.
.
R+U=200 . Substitute for R from above.
(2U+50)+U=200
3U+50=200
3U=150
U=50
.
R=2U+50=2(50)+50=100+50=150
.
ANSWER: Raheem gets Rs. 150 and Usman gets Rs. 50
.
8. Distribute an amount of Rs.500 among 2 brothers and 1 sister
such that,
a. sister gets twice as much as brothers do.
(Does this mean twice as much as each brother or twice as much as both?)
Using twice as much as each brother:
B=amount to each brother; S=amount to sister=2B
2B+S=500 . Substitute for S
2B+2B=500
4B=500
B=125
.
S=2B=2(125)=250
.
ANSWER: Each brother gets Rs. 125 and sister gets Rs. 250
b. each brother gets twice as much as the sister does.
B=amount to each brother; S=amount to sister
.
B=2S . Use this to substitute for B below.
.
2B+S=500 Substitute for B
2(2S)+S=500
4S+S=500
5S=500
S=100
.
B=2S=2(100)=200
.
ANSWER: Each brother gets Rs. 200 and sister gets Rs. 100